Suppose 3.4 g of zinc is treated with 9.0 g of pure hydrogen chloride. What is the maximum mass (in g) of hydrogen gas that can be produced? The balanced chemical equation is given below:

Zn + 2 HCl → ZnCl2 + H2

Express your answer to two significant figures.
1
Expert's answer
2020-05-26T14:21:17-0400

n(Zn)

"n=\frac{m}{M}=\frac{3.4}{65.38}=0.05"

n(HCL)

"n=\frac{m}{M}=\frac{9}{36.4609}=0.24"


decide further on zinc, in shortage

n(H2)=n(Zn)=0.05

m(H2)

"m=n\times M=0.05\times2.016=0.1" g